A delivery truck is transporting boxes of two sizes: large and small. The combined weight of a large box and a small box is 80 pounds. The truck is transporting 60 large boxes and 50 small boxes. If the truck is carrying a total of 4500 pounds in boxes, how much does each type of box weigh?
Answer by ikleyn(52781)

.
A delivery truck is transporting boxes of two sizes: large and small.
The combined weight of a large box and a small box is 80 pounds.
The truck is transporting 60 large boxes and 50 small boxes.
If the truck is carrying a total of 4500 pounds in boxes,
how much does each type of box weigh?
~~~~~~~~~~~~~~~


            It can be solved using arithmetic only,  without  Algebra.


Since the combined weight of one large box and one small box is 80 pounds,

50 large boxes and 50 small boxes  weight  50*80 = 4000 pounds, combined.


It means that remaining 60-50 = 10 large boxes weight 4500-4000 = 500  pounds.


Hence, one large box weights 500/10 = 50 pounds.


It means that each small box weights 80-50 = 30 pounds.

Solved  (mentally).